OpenStudy (cbreadbox):

solve x-3/n=b for x. Find any restrictions. I need help with this one. it's x-3 over n.

OpenStudy (tkhunny):

Notation Check - Order of Operations Which do you mean? \(x-\dfrac{3}{n} = b\;or\;\dfrac{x-3}{n} = b\) you have written the former, but you may mean the latter. Do NOT just clarify. Also show your work.

OpenStudy (texaschic101):

which one do you mean ?

OpenStudy (cbreadbox):

the second one, i mean

OpenStudy (cbreadbox):

so what I got... hold on..

OpenStudy (texaschic101):

lets see what tkhunny gets...he is smarter then me

OpenStudy (tkhunny):

You did not respond correctly. You seem to have missed the "show your work" part. Go!

OpenStudy (cbreadbox):

I was writng it out

OpenStudy (texaschic101):

I deleted

OpenStudy (tkhunny):

...and well you should write it out. Do it on here so we can see!

OpenStudy (cbreadbox):

ok

OpenStudy (tkhunny):

gtg @texaschic101 will take care of you.

OpenStudy (cbreadbox):

so I did x-3 over n times n=b times n *multiplied each term by n*

OpenStudy (cbreadbox):

then x-3=b*n *simplifies left hand side by cancelling out common factors*

OpenStudy (texaschic101):

correct, so far.....now add 3 to both sides

OpenStudy (texaschic101):

because you want x by itself

OpenStudy (cbreadbox):

x=3+bn??

OpenStudy (texaschic101):

yep...thats it

OpenStudy (cbreadbox):

ok thanks I needed to know it was correct.

OpenStudy (texaschic101):

it is correct :)

OpenStudy (cbreadbox):

what did it mean by "find any restrictions."?

OpenStudy (texaschic101):

I am not sure...don't know this one...sorry

OpenStudy (cbreadbox):

ok, thanks anyway!

OpenStudy (tkhunny):

n is in the denominator \(n \ne 0\)